OpenStudy (anonymous):

What is the perimeter of a triangle with the given side lengths? s1 = 2x + 1 cm s2 = 3x + 2 cm s3 = x + 3 cm P = x + cm

OpenStudy (anonymous):

The perimeter of a triangle with the side lengths s1 = 2x + 1 cm, s2 = 3x + 2 cm, s3 = x + 3 cm is: (2x + 1) + (3x + 2) + (x + 3) = 6x + 6 cm 6x+6cm is the answer

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np